# Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Streams

### v3

@@ Abstract
Particle Image Velocimetry (PIV) is a widely adopted non-invasive imaging technique that tracks the motion of tracer particles across image sequences to capture the velocity distribution of fluid flows. It is commonly employed to analyze complex flow structures and validate numerical simulations. This study explores the untapped potential of spike cameras—ultra-high-speed, high-dynamic-range vision sensors—in high-speed fluid velocimetry. We propose a deep learning framework, Spike Imaging Velocimetry (SIV), tailored for high-resolution fluid motion estimation. To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR). Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), comprising labeled samples from three representative scenarios in fluid dynamics. Our proposed method outperforms existing baselines on the PSSD dataset, demonstrating its effectiveness. The implementation details of our project, including code and configuration settings, are provided in the supplementary material.
Abstract
Particle Image Velocimetry (PIV) is a widely adopted non-invasive imaging technique that tracks the motion of tracer particles across image sequences to capture the velocity distribution of fluid flows. It is commonly employed to analyze complex flow structures and validate numerical simulations. This study explores the untapped potential of spike cameras—ultra-high-speed, high-dynamic-range vision sensors—in high-speed fluid velocimetry. We propose a deep learning framework, Spike Imaging Velocimetry (SIV), tailored for high-resolution fluid motion estimation. To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR). Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), comprising labeled samples from three representative scenarios in fluid dynamics. Our proposed method outperforms existing baselines on the PSSD dataset, demonstrating its effectiveness. The implementation details of our project, including code and configuration settings, are provided in the supplementary material.
@@ Introduction
The need for accurate and non-invasive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid dynamics research. This technique has been widely applied to critical topics such as studying the aerodynamic shapes of aircraft (Kompenhans et al. 2000), optimizing vehicle designs to reduce air resistance (Nakagawa et al. 2016), improving energy conversion efficiency in combustion processes (Chen, Li, and Zhu 2022), understanding the impact of blood flow dynamics on diseases (Raschi et al. 2012), and predicting the spread of water pollution (MoayeriKashani et al. 2016).
@@
Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetS (Cai et al. 2019b) and RAFT-PIV (Lagemann et al. 2021).
In PIV experiments or simulations, tracer particles (e.g., silver-coated particles) are introduced into the fluid to scatter optical signals. By estimating their displacements, PIV can infer the velocities of the fluid in which they are suspended. The success of optical flow networks (Dosovitskiy et al. 2015; Sun et al. 2018; Teed and Deng 2020) has inspired the development of learning-based PIV algorithms (Cai et al. 2019b, a; Lagemann et al. 2021). Although optical flow networks generally exhibit good generalization ability, the characteristics of fluid scenes cannot be ignored for further performance improvement. Turbulence is an irregular flow phenomenon occurring at high Reynolds numbers. When measuring the turbulent velocity field with PIV, this scenario is significantly different from macroscopic worlds, with the following key challenges:
(1) The fluid often moves rapidly within the field of view.
(2) The scene contains extensive empty regions without tracer particles, yet the network is required to estimate the full flow field.
(3) Unlike solids, the fluid is highly sparse and unstructured overall, despite the presence of vortices of different scales (called quasi-order structures).
These challenges often coexist in turbulence. Challenge (1) places high demands on the experimental hardware, typically requiring ultra-high-speed cameras. Challenge (2) requires addressing potential resolution loss in the network carefully. Challenge (3) introduces problems to the extraction of context features. To address these challenges, we propose improvements in both hardware and algorithm aspects.
In terms of hardware, we explore the application of spike cameras (a type of neuromorphic camera) for PIV. We argue that using spike cameras in fluid measurement tasks offers several advantages:
@@
High temporal resolution (HTR). Spike cameras can achieve ultra-high temporal resolution (20000Hz), reducing the interval $\Delta t$ between the source and target time, and providing rich temporal information for accurate predictions. This enables spike cameras to handle Challenge (1) effectively;
High dynamic range (HDR). When particles accumulate in large numbers, local regions may become overly bright, exceeding the dynamic range of conventional cameras. Spike cameras asynchronously accumulate photons at each pixel, inherently possessing a much larger dynamic range. This makes them more advantageous in scenarios with uneven brightness (such as inappropriate lighting or particles moving toward concentrated regions);
@@
Reduced experimental complexity. Spike cameras can continuously capture images at high frame rates with exceptionally low bandwidth and eliminate the need for synchronization controllers to modulate the camera and laser.
@@
We propose a specialized network, Spike Imaging Velocimetry (SIV), for estimating fluid velocity fields. To enhance network performance, we develop three novel modules tailored to the characteristics of fluid dynamics and spike cameras. These include the Detail-Preserving Hierarchical Transform (DPHT), which enhances feature extraction from spike streams while mitigating downsampling artifacts; the Graph Encoder (GE), which maps features into a graph structure for adaptive context aggregation; and the Multi-scale Velocity Refinement (MSVR), which addresses over-smoothing in predicted flow fields.
We also introduce a new dataset, Particle Scenes with Spike and Displacement (PSSD), designed for learning fluid motion from spike data. It covers three representative motion estimation scenarios: steady turbulence, high-speed flow, and high dynamic range conditions.
Extensive experiments demonstrate the superiority of SIV over both image-based and spike-based baselines. SIV achieves state-of-the-art performance across all evaluation settings. The effectiveness of each proposed module is further validated through ablation studies.
The main contributions of this work are summarized as follows:
1.
@@
We explore a feasible solution for applying spike cameras to fluid velocity measurement, opening new directions for PIV.
@@
The spike camera (Dong, Huang, and Tian 2021; Huang 2019) is a type of neuromorphic camera (NeurCam). Inspired by biological visual neural structures and mechanisms, NeurCams emulate the encoding principles of the primate retina. Spike cameras mimic the encoding principles of the central fovea in the retina, providing continuous visual information with ultra-high temporal resolutions, high dynamic range(HDR), and significantly reduced energy consumption and latency, making them well-suited for high-speed motion scenarios. Their pixels independently accumulate photons and emit a spike when the accumulated intensity exceeds a certain threshold. This process maps the incoming photon stream into the spike stream, which preserves spatiotemporal relationships and can be reconstructed into images. For a detailed explanation of the working mechanism, refer to (Dong, Huang, and Tian 2021). The application of spike cameras in optical flow estimation provides significant insights for this study. Although optical flow estimation is generally used to estimate the motion of macroscopic objects, its principles and objectives are aligned with those of PIV. The techniques employed in PIV for image registration and displacement estimation are essentially a variant of optical flow estimation. Spike-based optical flow estimation is still a relatively new field. (Hu et al. 2022) introduced a specialized neural network architecture for spike-based optical flow estimation. This structure can adaptively eliminate motion blur in spike streams based on prior motion information. A framework based on RAFT is proposed by (Zhao et al. 2022) to handle continuous spike streams. (Zhao et al. 2024) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network. (Chen, Yu, and Huang 2023) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by (Xia et al. 2024).
@@
Many traditional image-based PIV algorithms (Lourenco and Krothapalli 1995; Nobach and Honkanen 2005) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation windows. Displacement is then determined by finding the maximum correlation between the two interrogation windows, which is similar to the correlation method in optical flow estimation to some extent. However, traditional methods depend on finely tuned parameters and complex optimization techniques. With advancements in optical flow networks, various architectures (Dosovitskiy et al. 2015; Sun et al. 2018; Teed and Deng 2020) have been developed to predict dense velocity fields of macroscopic objects with high accuracy. Early PIV networks like PIV-DCNN (Lee, Yang, and Yin 2017) laid the groundwork for these developments. PIV-NetS (Cai et al. 2019b) is capable of generating full-resolution displacement outputs. (Cai et al. 2019a) adapt LiteFlowNet (Hui, Tang, and Loy 2018) to PIV. RAFT-PIV (Lagemann et al. 2021), derived from RAFT (Teed and Deng 2020), has shown exceptional performance on multiple datasets. (Yu et al. 2021) removes the computationally inefficient components of RAFT, such as the context encoder, to improve inference speed. Additionally, unsupervised learning frameworks for PIV(Zhang and Piggott 2020; Zhang et al. 2023a) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements. Other recent PIV networks include LIMA (Manickathan, Mucignat, and Lunati 2023), CC-FCN (Gao et al. 2021), En-FlowNetC (Ji, An, and Hui 2024), ARAFT-FlowNet (Han and Wang 2023), OFVNetS (Ji, Hui, and An 2024), and PIV-PWCNet (Zhang et al. 2023b). However, the performance gains of most PIV networks are largely attributed to advances in optical flow estimation architectures, rather than targeted enhancements tailored to the specific characteristics of fluid dynamics, and most implementations remain unavailable as open-source.
@@
The event camera (Lichtsteiner, Posch, and Delbruck 2008; Moeys et al. 2017; Posch, Matolin, and Wohlgenannt 2010; Huang, Guo, and Chen 2017) is another type of NeurCam, which is inspired by the peripheral retina (Rebecq et al. 2019). Unlike spike cameras, event cameras independently monitor the change in light intensity (increase or decrease) at each pixel and generate signals based on the local pixel information of dynamic changes. When the brightness change of a pixel exceeds a preset threshold, the camera records this change event (brightness increase as a positive event, brightness decrease as a negative event) and outputs its timestamp, position, and change value as data. However, due to the lack of absolute intensity values, event cameras struggle to reconstruct complete images, particularly in static regions and with fine textures. The application of event cameras to PIV has already been explored by (Willert and Klinner 2022) and (Willert 2023). Similar to the spike camera, the event camera offers higher temporal resolution, high dynamic range, low latency, and reduced experimental complexity.
@@
$A(\mathbf{x},t)=\int_{0}^{t}\alpha\cdot I(\mathbf{x},\tau)\bmod\theta,$ (1)
Where $A(\mathbf{x},t)$ represents the voltage of the accumulator at pixel $\mathbf{x}=(x,y)$ at time $\tau$, and $\theta$ is the threshold of the comparator. The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represents the temporal dimension. Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and brightness representation.
@@
Our proposed dataset, Particle Scenes with Spike and Displacement (PSSD), includes three sub-datasets: Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3). Each sub-dataset comprises 10,000 sample sequences of spike streams and the corresponding ground-truth turbulent velocity fields when $\Delta t=21$ (approximately 1.05 ms) and $\Delta t=11$ (approximately 0.55 ms). Additionally, we provide corresponding pairs of images for comparative analysis for image-based methods. We use the reference flow field to simulate the motion of the scene and construct a virtual spike camera to trigger the spike stream. The reference flow field data are taken from the high-precision numerical simulation database used for turbulence research, Johns Hopkins Turbulence Database (JHTDB) (Li et al. 2008). JHTDB employs Direct Numerical Simulation (DNS) to simulate turbulent flow fields. DNS directly solves the Navier-Stokes equations without any model assumptions or simplifications, enabling precise capture of all scales in turbulent flow. This method resolves the fluid motion in both space and time with high-resolution grids, accurately capturing the complex and multi-scale nature of turbulence. Additional details about the dataset can be found in our supplementary materials.
@@
$\mathcal{I}(\mathbf{p},t)\leftarrow\mathcal{I}(\mathbf{p},t+\Delta t).$ (2)
@@
In order to avoid the loss of particle signal information during feature extraction, we propose a multi-scale pyramidal architecture, named Detail-Preserving Hierarchical Transform (DPHT), which incorporates a 3D Detail-Preserving Downsampling module (as shown in Figure 3(a)).
3D Detail-Preserving downsampling. As shown in Figure 3(a), suppose the input of level $l$ is a feature vector $\mathbf{x}^{(l)}$, and the processing of $\mathbf{x}^{(l)}$ at this level is as follows:
@@
Here, $\mathscr{F}_{\mathrm{3C}}$ represents a 5×3×3 3D convolution, $\mathscr{F}_{\mathrm{3MP}}$ is a combination of a 3D max pooling operation and a 5×1×1 3D convolution, and $\mathscr{F}_{\mathrm{2C}}$ refers to a sequential stack of 2D convolutions. The output of the previous level serves as the input to the next level, i.e., $\mathbf{x}^{(l+1)}=\mathbf{x}^{(l)}_{\mathrm{out}}=\mathscr{D}^{(l)}(\mathbf{x}^{(l)})$.
Global temporal aggregation. In this step, the output feature vectors from each level are further refined by several convolutions (for detailed information, refer to the supplementary materials), and then concatenated together to form the final representation of the DPHT:
$\mathbf{R}=\mathscr{A}_{m}[\mathbf{Concat}(\mathbf{x}^{(l=1)}_{\mathrm{out}},\mathbf{x}^{(l=2)}_{\mathrm{out}},\mathbf{x}^{(l=3)}_{\mathrm{out}})$], (4)
$\mathscr{A}_{m}$ denotes convolution. After processing with DPHT, the signals of the source and target sub-streams are denoted as $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$, respectively.
@@
When the Reynolds number is high, fluid flow becomes turbulent. As energy cascades and dissipates, large vortices become unstable and gradually break into smaller ones until they vanish. This indicates that fluid motion is not purely chaotic, but rather exhibits an underlying topological structure. To capture the contextual information of the fluid, we propose a Graph Encoder (GE). It is a feature extraction network based on Graph Neural Networks (GNNs) and Convolutional Neural Networks (CNNs). By representing features as graph nodes and constructing a dynamic adjacency matrix, GE flexibly aggregates and updates information from neighboring nodes, enabling a larger receptive field than conventional CNNs. Graph Attention Networks (GATs) (Veličković et al. 2017), leveraging attention mechanisms, learn the relative contribution of each neighbor to a node’s new representation, thereby mitigating interference from particle-sparse regions. The serial combination of GNN and CNN enables adaptive global perception and local refinement, effectively capturing fluid-specific contextual features.
Graph projection. The first step is to map $\mathbf{R}_{s}$ extracted from DPHT from the general coordinate space to the node space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{s})$. The mapped context node representations, denoted as $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}\in\mathbb{R}^{C\times K}$, contain appearance features related to the shape and regional information of the scene. $C$ indicates channel number and $K$ is a hyper-parameter representing the number of nodes. In this paper, we set $K=128$. $\mathscr{P}_{R\to V}(\cdot)$ is a linear combination of feature vectors within the grid space, formulated by,
$v_{c,i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{s,j}).$ (5)
Here, $\mathscr{N}(\cdot)$ is the $L_{2}$-normalization function applied to the channel dimension of each node vector, and $z_{ij}$ denotes the projection weights computed by a CNN, i.e., $z=\mathscr{F}_{R\to V}(\mathbf{R}_{s})\in\mathbb{R}^{(H\times W)\times K}$.
Graph convolution layer. After obtaining the nodes $\mathbf{v}_{c}$ in the graph space, the dynamic adjacency matrix $\widetilde{\mathbf{A}}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ can be derived by calculating the similarity between all node vectors. The detailed definition of GAT can be found in the supplementary material. Here, we implicitly denote the operations of the two GAT modules as $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$. The updated node features are given by
@@
Graph re-projection. Our graph re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}=[z_{ij}]$ as inputs, generating a 2D feature map $\widetilde{\mathbf{R}}_{c}$. Since inverting the projection weight matrix is not feasible, we follow the approach in (Li and Gupta 2018; Luo et al. 2022) and compute the enhanced grid context features $\widetilde{\mathbf{R}}_{c}$ by re-weighting the vertex features $\mathbf{V_{c}^{\mathrm{out}}}$. This is given by,
$\widetilde{\mathbf{R}}_{c}=\mathscr{P}_{V\to R}(\mathbf{V_{c}^{\mathrm{out}}},\mathbf{Z})=\mathbf{Z}\mathbf{V_{c}^{\mathrm{out}}}^{T}.$ (7)
Thus, $\mathscr{P}_{V\to R}$ linearly interpolates 2D pixel features based on the node assignments of the 2D grid features, without additional parameters. Note that even if two pixels are assigned to the same node, they will have different features after re-projection. Therefore, the GCU may preserve the spatial details of the signal. Finally, these projection results can be integrated into a residual extractor network $\mathscr{R}$ for local refinement. Final context feature $\mathbf{F}_{c}$ given by,
@@
where $\alpha$ denotes a learnable parameter that is initialized as 0 and gradually performs a weighted sum.
@@
We adopt the Multi-scale Iterative Optimizer (MSIO) from DIP (Zheng et al. 2022) as the backbone for iterative residual flow estimation in SIV. Unlike RAFT’s full correlation volume construction, MSIO leverages Inverse PatchMatch to significantly reduce the computational cost introduced by large-scale feature maps (1/4 resolution of the full size). The detailed architecture of MSIO can be found in (Zheng et al. 2022); here, we denote it implicitly as $\mathscr{U}$. The output multi-scale velocity field is given by
$\{\mathbf{Vel}^{i}\}_{i=1}^{\text{iters}}=\mathscr{U}(\mathbf{F}_{c},\mathbf{F}_{m}),$ (9)
$\mathbf{F}_{m}$ denotes the motion features extracted by the Motion Encoder. iter denotes the number of iterations of GRU (Teed and Deng 2020). The detailed architecture of the Motion Encoder is provided in the supplementary material.

### v4

@@ Abstract
Particle Image Velocimetry (PIV) is a widely adopted non-invasive imaging technique that tracks the motion of tracer particles across image sequences to capture the velocity distribution of fluid flows. It is commonly employed to analyze complex flow structures and validate numerical simulations. This study explores the untapped potential of spike cameras—ultra-high-speed, high-dynamic-range vision sensors—in high-speed fluid velocimetry. We propose a deep learning framework, Spike Imaging Velocimetry (SIV), tailored for high-resolution fluid motion estimation. To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR). Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), which contains labeled samples from three representative fluid-dynamics scenarios: steady turbulence, high-speed flow, and high-dynamic-range conditions. Our proposed method outperforms existing baselines across all these scenarios, demonstrating its effectiveness.
@@ Introduction
Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetS (pivnets) and RAFT-PIV (raftpiv).
The need for accurate and non-invasive flow measurement techniques has led to the widespread use of Particle Image Velocimetry (PIV), a key diagnostic tool for fluid dynamics. PIV has been applied to critical areas such as studying aircraft aerodynamics (kompenhans2000particle), optimizing vehicle designs for reduced air resistance (nakagawa2016typical), improving energy efficiency in combustion processes (chen2022investigations), understanding blood flow dynamics in diseases (raschi2012cfd), and predicting water pollution spread (moayerikashani2016tracking). In PIV experiments, tracer particles scatter light within a fluid, and their displacements are used to infer the velocity field of the fluid. Recent advancements in optical flow networks (flownet; pwcnet; raft) have inspired the development of learning-based PIV algorithms (pivnets; pivliteflownet; raftpiv). While optical flow networks are highly generalizable, fluid scenes present unique challenges that limit further performance improvements. Turbulence, a chaotic flow phenomenon at high Reynolds numbers, introduces specific issues in velocity field estimation: (1) High-speed fluid movement within the field of view. (2) Sparse regions without tracer particles, leading to discrete signals while still requiring the estimation of a continuous flow field. (3) Unstructured patterns and shapes with small-scale vortices and quasi-ordered structures. These challenges coexist in turbulent flows, necessitating advanced experimental hardware and algorithmic solutions. To address these, we propose Spike Imaging Velocimetry (SIV). In terms of hardware, we explore the use of spike cameras (firstspike; spikepatent) for PIV. Spike cameras offer several advantages for fluid velocity measurement:
@@
- High Temporal Resolution (HTR): With a temporal resolution of 20,000 Hz, spike cameras significantly reduce the time interval $\Delta t$ between consecutive frames, providing rich temporal information that enables precise predictions. This addresses Challenge (1) effectively.
- High Dynamic Range (HDR): Spike cameras asynchronously accumulate photons at each pixel, providing a larger dynamic range, which is beneficial in scenarios with uneven lighting or high particle concentration. This helps mitigate limitations seen in conventional cameras, addressing Challenge (2).
@@
- Reduced Experimental Complexity: Spike cameras can continuously capture high-speed images with low bandwidth and do not require synchronization controllers, simplifying experimental setups.
@@
We propose a specialized network, Spike Imaging Velocimetry (SIV), for estimating fluid velocity fields. To enhance network performance, we develop three novel modules tailored to the characteristics of fluid dynamics and spike cameras. These include the Detail-Preserving Hierarchical Transform (DPHT), which enhances feature extraction from spike streams while mitigating downsampling artifacts; the Graph Encoder (GE), which maps features into a graph structure for adaptive context aggregation; and the Multi-scale Velocity Refinement (MSVR), which addresses over-smoothing in predicted flow fields. We also introduce a new dataset, Particle Scenes with Spike and Displacement (PSSD), designed for learning fluid motion from spike data. It covers three representative motion estimation scenarios: steady turbulence, high-speed flow, and high dynamic range conditions. Extensive experiments demonstrate the superiority of SIV over both image-based and spike-based baselines. SIV achieves state-of-the-art performance across all evaluation settings. The effectiveness of each proposed module is further validated through ablation studies. The contributions of this work are summarized as follows:
@@
1. We explore a feasible solution for applying spike cameras to fluid velocity measurement, opening new directions for PIV.
@@
The spike camera (firstspike; xiong2024real; su2024intensity) is a type of neuromorphic camera (NeurCam). Inspired by biological visual neural structures and mechanisms, NeurCams emulate the encoding principles of the primate retina. Spike cameras mimic the encoding principles of the central fovea in the retina, providing continuous visual information with ultra-high temporal resolutions, high dynamic range (HDR), and significantly reduced energy consumption and latency, making them well-suited for high-speed motion scenarios. Their pixels independently accumulate photons and emit a spike when the accumulated intensity exceeds a certain threshold. This process maps the incoming photon stream into the spike stream, which preserves spatiotemporal relationships and can be reconstructed into images. For a detailed explanation of the working mechanism, refer to (firstspike). The application of spike cameras in optical flow estimation provides significant insights for this study. Although optical flow estimation is generally used to estimate the motion of macroscopic objects, its principles and objectives are aligned with those of PIV. The techniques employed in PIV for image registration and displacement estimation are essentially a variant of optical flow estimation. Spike-based optical flow estimation is still a relatively new field. (hu2022optical) introduced a specialized neural network architecture for spike-based optical flow estimation. This structure can adaptively eliminate motion blur in spike streams based on prior motion information. A framework based on RAFT is proposed by (spike2flow) to handle continuous spike streams. (hist) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network. (chen2023self) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by (xia2024unsupervised).
@@
Many traditional image-based PIV algorithms (classicpiv1; classicpiv3) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation windows. Displacement is then determined by finding the maximum correlation between the two interrogation windows, which is similar to the correlation method in optical flow estimation to some extent. However, traditional methods depend on finely tuned parameters and complex optimization techniques. With advancements in optical flow networks, various architectures (flownet; pwcnet; raft) have been developed to predict dense velocity fields of macroscopic objects with high accuracy. Early PIV networks like PIV-DCNN (pivdcnn) laid the groundwork for these developments. PIV-NetS (pivnets) is capable of generating full-resolution displacement outputs. (pivliteflownet) adapt LiteFlowNet (liteflownet) to PIV. RAFT-PIV (raftpiv), derived from RAFT (raft), has shown exceptional performance on multiple datasets. (lightpivnet) removes the computationally inefficient components of RAFT, such as the context encoder, to improve inference speed. Additionally, unsupervised learning frameworks for PIV(unsupervisedpiv1; unsupervisedpiv2) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements. Other recent PIV networks include LIMA (manickathan2023lightweight), CC-FCN (gao2021robust), En-FlowNetC (ji2024cross), ARAFT-FlowNet (han2023attention), OFVNetS (ji2024effects), and PIV-PWCNet (zhang2023pyramidal). However, the performance gains of most PIV networks are largely attributed to advances in optical flow estimation architectures, rather than targeted enhancements tailored to the specific characteristics of fluid dynamics, and most implementations remain unavailable as open-source.
@@
The event camera (dvs; davis; atis; jiang2024edformer) is another type of NeurCam, which is inspired by the peripheral retina (event1). Unlike spike cameras, event cameras independently monitor the change in light intensity (increase or decrease) at each pixel and generate signals based on the local pixel information of dynamic changes. When the intensity change of a pixel exceeds a preset threshold, the camera records this change event (intensity increase as a positive event, intensity decrease as a negative event) and outputs its timestamp, position, and change value as data. The success of event cameras in optical flow estimation (lee2020spike-flownet; shiba2022secrets) has paved the way for their application to PIV. (EBIV) introduced the first event-camera PIV algorithm based on correlation and motion compensation. (eventpiv) combined an event camera with pulsed illumination, substantially reducing estimation latency. These studies are insightful and practical; however, their algorithmic novelty is limited, and robustness is partly constrained by the event camera’s lack of absolute intensity.
@@
$A(\mathbf{x},t)=\left(\int_{0}^{t}\alpha\cdot I(\mathbf{x},\tau)d\tau\right)\bmod\theta,$ (1)
Where $A(\mathbf{x},t)$ represents the voltage of the accumulator at pixel $\mathbf{x}=(x,y)$ at time $\tau$, and $\theta$ is the threshold of the comparator. The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represents the temporal dimension. Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and intensity representation.
@@
Training neural networks using supervised strategies requires data with ground truth to optimize model parameters. However, it is difficult to obtain accurate velocity fields from experimental PIV data. We have to generate a synthetic dataset for training, following (pivnets). Our proposed dataset, Particle Scenes with Spike and Displacement (PSSD), includes three sub-datasets: Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3). Each sub-dataset comprises 10,000 sample sequences of spike streams and the corresponding ground-truth turbulent velocity fields when $\Delta t=21$ (approximately 1.05 ms) and $\Delta t=11$ (approximately 0.55 ms). Additionally, we provide corresponding pairs of images for comparative analysis for image-based methods. We use the reference flow field to simulate the motion of the scene and construct a virtual spike camera to trigger the spike stream. Both the images and the spike stream in the dataset include added noise. The reference flow fields are obtained from the Johns Hopkins Turbulence Database (JHTDB) (JHTDB), a high-precision numerical simulation database widely used in turbulence research. JHTDB is based on direct numerical simulation (DNS) of the Navier–Stokes equations without any turbulence modeling, resolving all dynamically relevant scales in both space and time on fine grids. As a result, the simulated flow fields closely approximate real turbulent flows and can be regarded as ground truth in our study.
Figure 2: (a) Normal illumination scene (Problem 1, Problem 2); (b) High dynamic range scene (Problem 3).
@@
$\mathcal{I}(\mathbf{p},t)\leftarrow\mathcal{I}(\mathbf{p}+\mathbf{D}(\mathbf{p}),t+\Delta t).$ (2)
@@
To preserve particle signal information during feature extraction, we propose the Detail-Preserving Hierarchical Transform (DPHT), a multi-scale pyramidal architecture that integrates a 3D Detail-Preserving Downsampling module (see Figure 4(a)).
3D Detail-Preserving Downsampling. Given an input feature vector $\mathbf{x}^{(l)}$ at level $l$, the processing is performed as follows:
@@
where $\mathscr{F}_{\mathrm{3C}}$ is a 5×3×3 3D convolution, $\mathscr{F}_{\mathrm{3MP}}$ combines 3D max pooling and a 5×1×1 3D convolution, and $\mathscr{F}_{\mathrm{2C}}$ refers to a stack of 2D convolutions. The output of each level is passed to the next, i.e., $\mathbf{x}^{(l+1)}=\mathbf{x}^{(l)}_{\mathrm{out}}$.
Global Temporal Aggregation. The output feature vectors from each level are refined through additional convolutions and then concatenated to form the final DPHT representation:
$\mathbf{R}=\mathscr{A}{m}[\mathbf{Concat}(\mathbf{x}^{(l=1)}_{\mathrm{out}},\mathbf{x}^{(l=2)}_{\mathrm{out}},\mathbf{x}^{(l=3)}_{\mathrm{out}})$], (4)
where $\mathscr{A}_{m}$ denotes a convolution operation. The source and target sub-streams’ signals are represented as $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$, respectively.
@@
In turbulent flows with high Reynolds numbers, large vortices break into smaller ones, revealing an underlying topological structure in fluid motion. To capture this structure, we propose the Graph Encoder (GE), a feature extraction network based on Graph Neural Networks (GNNs) and Convolutional Neural Networks (CNNs). By representing features as graph nodes and constructing a dynamic adjacency matrix, GE aggregates and updates information from neighboring nodes, allowing for a larger receptive field than traditional CNNs. Graph Attention Networks (GATs) (GAT) leverage attention mechanisms to learn the relative contribution of each neighbor, reducing interference from particle-sparse regions. This combination of GNN and CNN enables adaptive global perception and local refinement, effectively capturing fluid-specific contextual features.
Graph Projection. The first step is to map $\mathbf{R}_{s}$ from the DPHT to the node space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{s})$. The context node representations, $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}\in\mathbb{R}^{C\times K}$, contain appearance features related to the scene’s shape and regional information, where $C=128$ is the channel number and $K$ is the number of nodes (set to $K=128$ in this paper). The projection is formulated as:
$v_{c}^{i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{s,j}),$ (5)
where $\mathscr{N}(\cdot)$ is the $L_{2}$ normalization, and $z_{ij}$ denotes the normalized projection weights computed by a $1\times 1$ convolution and a softmax function.
Graph Convolution Layer. After obtaining the nodes $\mathbf{v}_{c}$, the dynamic adjacency matrix $\widetilde{\mathbf{A}}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ is computed based on the similarity between node vectors. The operations of two GAT modules, $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$, update the node features as follows:
@@
Graph Re-projection. The re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}$, producing a 2D feature map $\widetilde{\mathbf{R}}c$:
$\widetilde{\mathbf{R}}_{c}=\mathscr{P}_{V\to R}(\mathbf{V_{c}^{\mathrm{out}}},\mathbf{Z})=\mathbf{Z}\left(\mathbf{V_{c}^{\mathrm{out}}}\right)^{T}.$ (7)
This operation preserves spatial details, as even if two pixels share the same node, they will have distinct features. Finally, these results are integrated into a residual extractor network $\mathscr{R}$ for local refinement:
@@
where $\alpha$ is a learnable parameter initialized to 0, gradually refining the feature sum.
@@
We adopt the Multi-scale Iterative Optimizer (MSIO) from DIP (dip) as the backbone for iterative residual flow estimation in SIV. Unlike RAFT’s full correlation volume construction, MSIO leverages Inverse PatchMatch to significantly reduce the computational cost introduced by large-scale feature maps (1/4 resolution of the full size). The detailed architecture of MSIO can be found in (dip); here, we denote it implicitly as $\mathscr{U}$. The output multi-scale velocity field is given by
$\{\mathbf{u}_{i}\}_{i=1}^{N}=\mathscr{U}(\mathbf{F}_{c},\mathbf{F}_{m}),$ (9)
where $\mathbf{F}_{m}$ denotes motion features extracted by the Motion Encoder, and $N$ refers to the number of GRU iterations (raft).